OpenStudy (anonymous):

how to get the value of x in the equation log(2x+7)-log(x-1)=log 5

OpenStudy (anonymous):

Properties of logarithms!!! :D \[\large \log \frac{M}{N}=\log\ M - \log\ N ; \ \ \ \ \ \ \ N \neq 0\]

Parth (parthkohli):

Also,\[\large \log N = \log M \iff M = N\]:D

OpenStudy (anonymous):

thank you